OpenStudy (anonymous):

Write a polynomial function with rational coefficients so that p(x) = 0 has the given roots. 5, 2i

OpenStudy (sasogeek):

please don't spam your questions bro :) someone will come to your aid :)

OpenStudy (sasogeek):

you can ask it once and then ask for help in the chat and paste the link there ;) that's better and we'll all be happy :D

OpenStudy (anonymous):

where's the chat?

OpenStudy (sasogeek):

it's down the page there and it should be an orange color at the moment if you haven't clicked it yet... "Mathematics"

OpenStudy (sasogeek):

in any case if a polynomial has roots 5 and 2i, it means that you should have something like (x-5)(x-2i)=0 when you factorize it, so you expand that and you should have the polynomial function :)

OpenStudy (anonymous):

How would I expand it?

OpenStudy (sasogeek):

i will be there in a while but jorgeeeee here's how to expand it (x-5)(x-2i)=0 x(x-2i)-5(x-2i)=0 then use the distributive property to expand it :) let me know if you can't do that too, i'll help you with it :)

OpenStudy (anonymous):

x(x-2i)-5(x-2i)=0 so, that would be (x-5)(x-2i)^2 ?

OpenStudy (sasogeek):

nooo lol ok i'll teach you \(\huge a(b+c)=ab+ac \) so with what we have here, you'll have something like this \[\huge (x-5)(x-2i)=0 \]\[\huge x(x-2i)-5(x-2i)=0 \] then we'll apply the first rule here \[\huge (x^2-x(2i))-(5x-5(2i))=0 \]\[\huge x^2-2xi-5x+10i=0 \] do you understand it up to this point?

OpenStudy (anonymous):

What rule did you exactly apply? you're multiplying?

OpenStudy (sasogeek):

i'm multiplying yes but this is the distributive property :)

OpenStudy (anonymous):

So you always set it up the opposite way like (x+ or - and the number) then use the second one twice?

OpenStudy (sasogeek):